OpenStudy (anonymous):

Prove that: ((Sin x)^2 + 2 Cos X -1)/((Sin X)^2 + 3Cos X -3) = ((Cos X)^2 + Cos X)/-(Sin X)^2

OpenStudy (anonymous):

sometimes it helps to have sin x = y/r and cos x =x/r IDK might help

jhonyy9 (jhonyy9):

do you know how can you writing 1 like relation of sin^2 x and cos^2 x ?

OpenStudy (anonymous):

Yes Sin^2 X

jhonyy9 (jhonyy9):

do you know this formula sin^2(x) +cos^2(x) =1

OpenStudy (anonymous):

Yes but there's only so much you can do with that formula

jhonyy9 (jhonyy9):

so how do you think (sin x)^2 not equal sin^2(x) ?

OpenStudy (anonymous):

thats the same expression written differently

jhonyy9 (jhonyy9):

yes right so than check it in your exercise please what you will can rewriting there if you know thie what i have wrote there now

jhonyy9 (jhonyy9):

if you know this what ...

OpenStudy (anonymous):

\[(\sin ^{2} \theta+ 2\cos \theta -1)/(\sin ^{2} \theta + 3\cos \theta -3) = (\cos ^{2} \theta+\cos \theta)/-\sin ^{2}\theta\]

jhonyy9 (jhonyy9):

and ... ? so there you not have used thie relation of sin^2(x) +cos^2(x) in place of 1 ,why ?

jhonyy9 (jhonyy9):

this relation of ---

OpenStudy (anonymous):

I'm having trouble understand what you're trying to say

jhonyy9 (jhonyy9):

so if you know that 1=sin^2(x) +cos^2(x) so than why you not write inside your equation in place of 1 this sin^2(x)+cos^2(x) ? - so if you will write this there will can seeing what will get inside parantheses

jhonyy9 (jhonyy9):

come on i like make you understanding all this proof because will be very easy

jhonyy9 (jhonyy9):

little help (sin^2(x) +2cos x -1) so if there you write 1=sin^2(x)+cos^2(x) than what will get ? sin^2(x) +2cos x -((sin^2(x) +cos^2(x))) = sin^2(x) +2cos x -sin^2(x)-cos^2(x) = =2cos x -cos^2(x) - so yes ?

jhonyy9 (jhonyy9):

so in denominator there is 3cos x -3 - what can you factorizing out ? and what will get hence ?

jhonyy9 (jhonyy9):

can you understanding this all or ... difficile ?

OpenStudy (anonymous):

Oui je comprends

jhonyy9 (jhonyy9):

c'est bien mon amie - so if i have wrote it right ,correct ?

jhonyy9 (jhonyy9):

bon,nous pouvons continue ,je pense ,oui ?

jhonyy9 (jhonyy9):

ou apré cette moment tu peux continue seulment ?

OpenStudy (anonymous):

J'ai 1-Cos2^(X) + 3(Cos X -1) pour le denominateur

jhonyy9 (jhonyy9):

sorry i think my farnche not is correct because i have learned it in my schooles in 1987 was the last year

jhonyy9 (jhonyy9):

oh this sign that you have understood it all my words correct sure ?

OpenStudy (anonymous):

Yeah I used the trig identities

jhonyy9 (jhonyy9):

ok nice thank you for this courage for franche language so firstly check it please in numerator what will remaining there ?

OpenStudy (anonymous):

(2 Cos X - Cos^2 X)/(Sin^2 X + 3(Cos -1))

jhonyy9 (jhonyy9):

yes and now make this what i have maked in numerator write 1 in function of sin and cos - but inside parantheses you missed one x

jhonyy9 (jhonyy9):

je m'apelle André et tu ?

jhonyy9 (jhonyy9):

ok we make different steps - look if you know that 1=sin^2(x) +cos^2(x) what can you writing in denominator sin^2(x)= ... ?

jhonyy9 (jhonyy9):

- so for you can understanding more easy we need rewriting all terms of sin(x) in function of cos(x)

OpenStudy (anonymous):

yeah Sin^2(X) becomes 1-Cos2^x

jhonyy9 (jhonyy9):

yes than so what will get in denominator ?

OpenStudy (anonymous):

1- Cos2^X + 3(Cos X -1)

OpenStudy (anonymous):

I meant 1- Cos^2 X + 3(Cos X -1)

jhonyy9 (jhonyy9):

yes right but this sign of exponent you not write in correct place

jhonyy9 (jhonyy9):

oh y

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

and now can you rewriting all this exercise just using cos x so without sin x ?

jhonyy9 (jhonyy9):

sorry but my internet conection with server is very very slowly

OpenStudy (anonymous):

You can rewrite all Sin^2 X with 1- Cos^2 X but the Cos X is the problem

jhonyy9 (jhonyy9):

not is problem you need rwriting your exercise just without sin x ok ?

jhonyy9 (jhonyy9):

how will be ?

jhonyy9 (jhonyy9):

so you dont like tel me what is your really name ?

OpenStudy (anonymous):

1- Cos2^X + 3(Cos X -1) this is the denominator

OpenStudy (anonymous):

it becomes (2 Cos X - Cos^2 X)/(1- Cos2^X + 3(Cos X -1))

jhonyy9 (jhonyy9):

what we know that 2cos(x) -cos^2(x) cos^2(x) +cos(x) --------------------- = --------------- 1-cos^2(x) +3(cosx -1) (-(1-cos^2(x))) - yes ?

OpenStudy (anonymous):

2cos(x) -cos^2(x) cos^2(x) +cos(x) --------------------- = --------------- 1-cos^2(x) +3(cosx -1) (cos^2(x)-1)

jhonyy9 (jhonyy9):

yes is correct - so but now how do you like your exercise ,is more easy look because there are just cos of x ,yes ? - and in this case how do you think what can we make for solving this equation ?

OpenStudy (anonymous):

The question here is to prove one side is equal to the other side, I don't think solving for Cos of X is neccessary

jhonyy9 (jhonyy9):

ok is right but for you can makeing the calcules more easy you need note the cosx =y - ok ?

jhonyy9 (jhonyy9):

so than what will get ,how will be your exercise ?

jhonyy9 (jhonyy9):

2y-y^2 ------------- = 1-y^2 +3(y-1) -can you continue ?

jhonyy9 (jhonyy9):

y^2 +y = -------- y^2 -1 can you continue ?

jhonyy9 (jhonyy9):

come on we will ending imediatly !!!

OpenStudy (anonymous):

right side becomes y/(y-1)

jhonyy9 (jhonyy9):

so there are 2y-y^2 y^2 +y -------------- = ------- 1-y^2 +3(y-1) y^2 -1 yes you are right ,so the left parte ?

OpenStudy (anonymous):

I see how it goes

jhonyy9 (jhonyy9):

can you calcule till the end please ?

OpenStudy (anonymous):

y(2-y) -------------- (y-1)(2-y)

OpenStudy (anonymous):

left side = right side

jhonyy9 (jhonyy9):

ok correct and now you wann tel what is your really name and was my pleasure that i have helped to you in this nice proof hope so much that is ,that was understandably sure

OpenStudy (anonymous):

My name is Wulusi, Thanks!

jhonyy9 (jhonyy9):

thank you and you are really franche ?

OpenStudy (anonymous):

I studied in Quebec

jhonyy9 (jhonyy9):

in Canada ,ok and your birth country is Canada too ?

OpenStudy (anonymous):

Yeah I'm Canadian

jhonyy9 (jhonyy9):

thank you so my wife ask me what is the clock there now ?

OpenStudy (anonymous):

2:22 am

jhonyy9 (jhonyy9):

so the proof was understandably ?

OpenStudy (anonymous):

yeah I understand it

jhonyy9 (jhonyy9):

2:22 am what sign that there is saturday too like here in Budapest ,country Hungary but here is now 8:24

OpenStudy (anonymous):

ahhh good morning!

jhonyy9 (jhonyy9):

thank you and for you sweet dreams so if you will sleep after this now

OpenStudy (anonymous):

yeah thank you too!

jhonyy9 (jhonyy9):

pardon for my question but you have not name like Robert,or Michel or ... , or this name Wulusi is the same name or this is your family name ?

OpenStudy (anonymous):

Yeah that is my name, my parents are From China

jhonyy9 (jhonyy9):

ok was my pleasure good luck bye aurevoir - so if this is correct ?

jhonyy9 (jhonyy9):

and do you know my country ? or the capital city Budapest ? - here are so much people from China too - and you are girl or boy ?

OpenStudy (anonymous):

Hungary? I'm a boy

jhonyy9 (jhonyy9):

ok thank s ,yes Hungary ,in midle of Europe

jhonyy9 (jhonyy9):

and on what class you learn now ?

OpenStudy (anonymous):

I'm just doing this math question to help my friend, I'm in biotechnology

jhonyy9 (jhonyy9):

ok and hence how many years ,how old are you please ? i am 43

jhonyy9 (jhonyy9):

and my hobby is math i dont learn now

jhonyy9 (jhonyy9):

just i like help people - so if i can

OpenStudy (anonymous):

I'm 23 I'm not really good at math haha

jhonyy9 (jhonyy9):

ok thank s have a nice day and good luck byebye